OpenStudy (anonymous):

Use the distributive property to simply: 30x^2y^2+55xy --------------- over 5xy

OpenStudy (anonymous):

i am in the middle of a quiz and am stuck on this.

OpenStudy (pfenn1):

\[\frac{30x^2y^2+55xy}{5xy}=\frac{30x^2y^2}{5xy}+\frac {55xy}{5xy}\]Can you further simplify this?

OpenStudy (anonymous):

ya, is it gonna be 11/1?

OpenStudy (pfenn1):

No, simplify each term.

OpenStudy (anonymous):

so the 1st term would simplify to: 6/1xy right?

OpenStudy (anonymous):

noooooooooo dont leave meee!

OpenStudy (anonymous):

so the answer is 6xy + 11.

OpenStudy (pfenn1):

Correct!

OpenStudy (anonymous):

yaaaaaahhoooooooo! thanx

OpenStudy (pfenn1):

you're welcome

OpenStudy (anonymous):

see ya. in like a couple min on my next question.

OpenStudy (pfenn1):

ok, I'll try to keep an eye out for it.